OpenStudy (anonymous):

A cylinder has its height doubled and its radius cut to one third. What is the ratio of the volumes of the modified cylinder to the original cylinder?

OpenStudy (anonymous):

The initial cylinder's volume V(i) = h(i) * (pi * r(i)^2) right? The new volume V(n) = h(n) * (pi * r(n)^2). Knowing that r(i)/3 = r(n) and 2h(i) = h(n), can you see where to go from here?